OpenStudy (anonymous):

(-3x^3-3x^2+2x-4) / (x-1/3). Determine the remainder for the following using the remainder theorm. The remainder is different from 0 and is equal to??

jimthompson5910 (jim_thompson5910):

The remainder theorem is If P(x) is divided by x-k, then the remainder r is r = P(k)

jimthompson5910 (jim_thompson5910):

In this case, P(x) = -3x^3-3x^2+2x-4 and k = 1/3

OpenStudy (anonymous):

ok so how would i work this. I've worked it once and it was wrong.

jimthompson5910 (jim_thompson5910):

P(x) = -3x^3-3x^2+2x-4 P(1/3) = -3(1/3)^3-3(1/3)^2+2(1/3)-4 Keep going

OpenStudy (anonymous):

ok one second let me work it

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

i got -34/9

OpenStudy (anonymous):

would that be the end result?

jimthompson5910 (jim_thompson5910):

-34/9 is the remainder, so you are correct

OpenStudy (anonymous):

so that is the answer?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

i want to ask another one u going to be here a second? did u check my work on that one?

jimthompson5910 (jim_thompson5910):

yes, -34/9 is the correct answer

OpenStudy (anonymous):

It says to type it in as a simplifed fraction so would that be simplifed

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

yes it's fully simplified, it's the final answer

OpenStudy (anonymous):

i'm going to post another one will u help me on it?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

I posted it . :)